# Generative Diffusion Models for High Dimensional Channel Estimation

### v1

@@ I Introduction
As the most popular generative model, DMs’ proficiency in modeling intricate distributions and furnishing robust prior information underscores its promise in tackling inverse problems such as channel estimation. A MIMO channel estimator utilizing score-based generative models (SGMs), closely associated with DMs, was proposed in [25], achieving state-of-the-art estimation accuracy and robust out-of-distribution performance. However, the practical value of this scheme is hindered by the huge number of inference steps and network parameters involved, hardly satisfying the latency requirement of channel estimation while leading to a substantial storage burden. To address the challenge of real-time implementation, the authors of [26] develop a low-complexity DM-based channel estimator using a lightweight CNN, which learns the channel prior within the angular domain. However, the proposed method necessitates orthogonal pilot sequences to conduct an LS estimate as initialization, leading to substantial overhead in high-dimensional MIMO systems. Moreover, applying these DM-based estimators [25, 26] to quantized MIMO receivers with low-resolution ADCs remains an open issue. Additionally, existing works assume the availability of a large dataset containing clean channel samples for training DMs. However, collecting such a dataset can be difficult or even impossible for realistic air interfaces.
@@ II-A MIMO Channel Estimation
where $\mathbf{H}\in\mathbb{C}^{N_{\rm r}\times N_{\rm t}}$ represents the MIMO channel matrix to be estimated, and $\mathbf{n}_{k}\sim\mathcal{CN}(\mathbf{0},2\sigma_{n}^{2}\mathbf{I})$ denotes the additive white Gaussian noise (AWGN) vector with variance $\sigma_{n}^{2}$ per real/imaginary component. The MIMO channels are assumed to be quasi-static, remaining constant within the $N_{\rm p}$ pilot transmissions. Therefore, the received signal during this training period can be written in the matrix form $\mathbf{Y}\in\mathbb{C}^{N_{\rm r}\times N_{\rm p}}$ as
@@
where $\mathbf{P}=[\mathbf{p}_{1},\ldots,\mathbf{p}_{N_{\rm p}}]$ and $\mathbf{N}=[\mathbf{n}_{1},\ldots,\mathbf{n}_{N_{\rm p}}]$ contain the stacked pilot symbols and noise vectors, respectively. For simplicity, the pilot symbols are supposed to be equiprobably drawn from the quadrature phase shift keying (QPSK) lattice, i.e., {±$\{\pm 1/\sqrt{2}\pm j/\sqrt{2}\}$, fulfilling the unit power constraint.
@@
The objective of channel estimation is to estimate $\mathbf{h}$ given the received pilot measurements ${\mathbf{y}}$ and the knowledge of the pilot $\mathbf{P}$. Let $\alpha=N_{\rm p}/N_{\rm t}$ denote the pilot density. When $\alpha<1$, channel estimation becomes an under-determined linear inverse problem, requiring strong priors for regularization.
@@ III-A2 Derivation of the Posterior Inference Algorithm
To resolve the challenge in handling $\nabla_{\mathbf{h}_{t}}\log p_{t}(\mathbf{y}|\mathbf{h}_{t})$, we resort to approximations using an uninformative prior assumption inspired by [34]. Specifically, by assuming $p_{0}(\mathbf{h}_{0})$ is uninformative with respect to $p(\mathbf{h}_{t}|\mathbf{h}_{0})$, we have $p(\mathbf{h}_{0}|\mathbf{h}_{t})\propto p(\mathbf{h}_{t}|\mathbf{h}_{0})p_{0}(\mathbf{h}_{0})\approx p(\mathbf{h}_{t}|\mathbf{h}_{0})$. This assumption is asymptotically precise when $t\to 0$, as shown in the exemplified verification in [34, Appendix A]. To see why it works, recall that from (8) we have
@@ III-C Learning from Noisy Channel Realizations
Collectively, the proposed method adds noise to the MMSE-denoised channel samples obtained by SURE denoising and learns the corresponding generative prior based on DM’s training objective. The SURE denoising and DM’s training can be unified by Tweedie’s formula [47], enabling both parts to share the same network structure $\boldsymbol{\epsilon}_{\boldsymbol{\theta}}$ while using different parameters, ${\boldsymbol{\theta}}_{1}$ and ${\boldsymbol{\theta}}_{2}$. The block diagram of the proposed training scheme, termed SURE-DM, is presented in Fig. 2.
@@ IV-A Experimental Details
We adopt the QuaDRiGa toolbox [48] as the channel simulator to generate training, testing, and validation data. We consider an urban macro-cell line-of-sight (LOS) scenario according to the 3rd generation partnership project (3GPP) 38.901 technical report [49]. The carrier frequency is set to 40 GHz. ULAs with “3GPP-3D” antennas featuring half-wavelength spacing are utilized at both the base station (BS) and the multi-antenna user terminal (UT). Downlink transmission is assumed, where the BS’s antenna array is placed at a height of 25 m. The UT randomly resides in a $120^{\circ}$ cell sector with a radius of 500 m and a guard distance of 35 m to the BS. The number of antennas at the BS and UT is set as $N_{\rm t}=64$ and $N_{\rm r}=16$, respectively. A total of $D=100,000$ channel samples are generated as the training dataset. Two distinct setups are considered, i.e., noiseless and noisy datasets. For the noiseless case (Sections IV-B and IV-C), a simulated dataset =$\mathcal{H}=\{\mathbf{h}^{(i)}\}_{i=1}^{D}$ containing ground truth samples generated via QuaDRiGa is employed; for the noisy case (Section IV-D), the simulated dataset $\mathcal{H}$ serves as a benchmark for deriving normalized LMMSE estimates as given in (32), constituting a noisy dataset ¯=$\bar{\mathcal{H}}=\{\bar{\mathbf{h}}^{(i)}\}_{i=1}^{D}$. Furthermore, a validation dataset containing 100 channel realizations is utilized for hyperparameter tuning.
During training, the number of epochs is set to 500. The batch size is configured as 128. The Adam optimizer with the learning rate fixed as $10^{-4}$ is utilized. For the noise schedule $\{\beta_{t}\}_{t=1}^{T}$ of DM, we employ the linear schedule prescribed in [50, Table 1] for different numbers of time steps $T$. For the backbone CNN of the denoising network, the maximum channel size is set as $S_{\max}=64$, leading to the channel sizes of different convolutional layers as shown in Fig. 12. Details of the network are illustrated in Appendix A. The gradient scale $s$ of the proposed approach is selected by grid search using the validation dataset. During testing, we generate 100 channel realizations with UT locations that differ from the training dataset to assess the estimation normalized MSE (NMSE) performance defined as
@@
The pilot symbols are randomly drawn from the QPSK lattice unless noted otherwise.
@@
Hyperparameters of these baselines are meticulously tuned based on the validation dataset. Note that for LDAMP and VAE, a distinct model is trained for each value of SNR, $\alpha$, or quantization bit count to maximize their performance due to their supervised learning nature.
@@ IV-B Full-Resolution Channel Estimation
We first evaluate the estimation performance of the proposed DM-based method in the full-resolution receiver. Fig. 3 presents the NMSE as a function of the pilot density $\alpha$ with the SNR fixed as 30 dB. We consider three distinct configurations for the proposed method. The former two are equipped with $T=100$ and $T=1000$, respectively. The last one, also equipped with $T=100$ time steps, utilizes an enhancing strategy to refine the estimate by iterating the update rule 3 rounds at each time index $t\in[0,T/2]$, i.e., the latter half of the reverse process. This strategy, termed “DM T100-Enhanced”, is motivated by the observation that there exists a refinement stage as $t$ approaches 0 where additional efforts can significantly refine the details of the estimation [53]. As shown in Fig. 4(a), the “DM T100-Enhanced” variant outperforms the vanilla versions with both $T=100$ and $T=1000$, particularly when $\alpha$ is small. Similar phenomena are noted in the low SNR regime, showcasing the superiority of this variant. Therefore, this variant is maintained for the remaining experiments. Moreover, the figure shows that the NMSE improvement within the range $\alpha\in(0.4,1]$ is relatively gradual compared to the rapid decrease in estimation error within the range $\alpha\in[0.2,0.4]$. This result reveals that the prior incorporated in the DM can provide sufficient information for approaching the recovery performance achieved in a fully determined inverse problem.
@@
Fig. 4 compares the NMSE and latency of the proposed method with baselines when the pilot density is $\alpha=1$. As shown in Fig. 4(a), DM outperforms all the compared schemes by a wide margin in estimation accuracy except for the high-complexity SGM. Due to the absence of true covariance information, the LMMSE estimator only uses an estimated sample covariance, hence lagging behind the DM that can capture deep-level propagation features. Fig. 4(b) presents the online estimation latency per channel realization across the compared methods. The latency is evaluated on a machine with an NVIDIA Tesla V100 GPU and an Intel Xeon E5-4627 CPU. The figure demonstrates that the proposed method achieves notable latency reduction compared to the SGM, making it one of the most efficient methods among those compared. This advantage can be attributed to the lightweight network architecture, substantially reduced time steps, and low computational overhead at each step. Considering the comparable performance to SGM as depicted in Fig. 4(a), our method can be viewed as a favorable option for wireless channel estimation featuring real-time implementation.
@@
To verify the robustness of the proposed deep generative prior-based method, we consider a mismatched channel setting in Fig. 7. Specifically, we deploy the DM trained using LOS channel samples in a non-line-of-sight (NLOS) scenario (with other configurations remaining unchanged) for channel estimation without any retraining or adaptation. We also present the NMSE of benchmark models trained with NLOS channel samples. As shown in the figure, the performance loss of the proposed method in mismatched setups is minor compared to that of the LMMSE and akin to the SGM baseline. This can be attributed to the powerful distribution coverage capability of generative DMs [17].
@@ IV-C Low-Resolution Channel Estimation
Figure 10: Performance with various ADC resolutions when $\alpha=0.6$.
@@
IV-D Effectiveness of SURE-DM

### v2

@@ I Introduction
As the most popular generative model, DMs’ proficiency in modeling intricate distributions and furnishing robust prior information underscores its promise in tackling inverse problems such as channel estimation. A MIMO channel estimator utilizing score-based generative models (SGMs), closely associated with DMs, was proposed in [25], achieving state-of-the-art estimation accuracy and robust out-of-distribution performance. However, the practical value of this scheme is hindered by the huge number of inference steps and network parameters involved, hardly satisfying the latency requirement of channel estimation while leading to a substantial storage burden. To address the challenge of real-time implementation, the authors of [26] develop a low-complexity DM-based channel estimator using a lightweight CNN, which learns the channel prior within the angular domain. However, the proposed method necessitates orthogonal pilot sequences and full pilot measurements to conduct an LS estimate as initialization, leading to substantial overhead in high-dimensional MIMO systems. Moreover, applying these DM-based estimators [25, 26] to quantized MIMO receivers with low-resolution ADCs remains an open issue. Additionally, existing works assume the availability of a large dataset containing clean channel samples for training DMs. However, collecting such a dataset can be difficult or even impossible for realistic air interfaces.
@@ II-A MIMO Channel Estimation
where $\mathbf{H}\in\mathbb{C}^{N_{\rm r}\times N_{\rm t}}$ represents the MIMO channel matrix to be estimated, and $\mathbf{n}_{k}\sim\mathcal{CN}(\mathbf{0},2\sigma_{n}^{2}\mathbf{I})$ denotes the additive white Gaussian noise (AWGN) vector with variance $\sigma_{n}^{2}$ per real/imaginary component. The MIMO channels are assumed to be quasi-static, remaining constant within the $N_{\rm p}$ pilot transmissions. Therefore, the received signal during this training period can be written in the matrix form $\mathbf{Y}\in\mathbb{C}^{N_{\rm r}\times N_{\rm p}}$ as [^1] Although we assume a fully digital architecture in the system model, the proposed method can be readily extended to massive MIMO systems with hybrid analog-digital architectures.
@@
where $\mathbf{P}=[\mathbf{p}_{1},\ldots,\mathbf{p}_{N_{\rm p}}]$ and $\mathbf{N}=[\mathbf{n}_{1},\ldots,\mathbf{n}_{N_{\rm p}}]$ contain the stacked pilot symbols and noise vectors, respectively. For simplicity, the pilot symbols are supposed to be equiprobably drawn from the quadrature phase shift keying (QPSK) lattice, i.e., {±$\{\pm 1/\sqrt{2}\pm{\rm j}/\sqrt{2}\}$, fulfilling the unit power constraint.
@@
The objective of channel estimation is to estimate $\mathbf{h}$ given the received pilot measurements ${\mathbf{y}}$ and the knowledge of the pilot $\mathbf{P}$. Let $\alpha=N_{\rm p}/N_{\rm t}$ denote the pilot density. When $\alpha<1$, channel estimation becomes an under-determined linear inverse problem, requiring strong priors for regularization. [^2] In addition to leveraging strong prior information, such as deep generative priors, to reduce pilot overhead, deep learning-based channel extrapolation has also emerged as an effective alternative strategy [29]. However, as this paper focuses on a fundamentally different approach, the detailed mechanisms of channel extrapolation and its comparison with the proposed method are beyond the scope of this work.
@@ III-A2 Derivation of the Posterior Inference Algorithm
To resolve the challenge in handling $\nabla_{\mathbf{h}_{t}}\log p_{t}(\mathbf{y}|\mathbf{h}_{t})$, we resort to approximations using an uninformative prior assumption inspired by [35]. Specifically, by assuming $p_{0}(\mathbf{h}_{0})$ is uninformative with respect to $p(\mathbf{h}_{t}|\mathbf{h}_{0})$, we have $p(\mathbf{h}_{0}|\mathbf{h}_{t})\propto p(\mathbf{h}_{t}|\mathbf{h}_{0})p_{0}(\mathbf{h}_{0})\approx p(\mathbf{h}_{t}|\mathbf{h}_{0})$. This assumption is asymptotically precise when $t\to 0$, as shown in the exemplified verification in [35, Appendix A]. To see why it works, recall that from (8) we have
@@ III-C Learning from Noisy Channel Realizations
Collectively, the proposed method adds noise to the MMSE-denoised channel samples obtained by SURE denoising and learns the corresponding generative prior based on DM’s training objective. The SURE denoising and DM’s training can be unified by Tweedie’s formula [48], enabling both parts to share the same network structure $\boldsymbol{\epsilon}_{\boldsymbol{\theta}}$ while using different parameters, ${\boldsymbol{\theta}}_{1}$ and ${\boldsymbol{\theta}}_{2}$. We adopt separate weights for the SURE and DM due to the fundamentally different denoising objectives of the two components, with SURE targeting a single noise level and DM adapting to multiple noise levels. The block diagram of the proposed training scheme, termed SURE-DM, is presented in Fig. 2.
@@ IV-A Experimental Details
We adopt the QuaDRiGa toolbox [49] as the channel simulator to generate training, testing, and validation data. We consider an urban macro-cell line-of-sight (LOS) scenario according to the 3rd generation partnership project (3GPP) 38.901 technical report [52]. The carrier frequency is set to 40 GHz. ULAs with “3GPP-3D” antennas featuring half-wavelength spacing are utilized at both the base station (BS) and the multi-antenna user terminal (UT). Downlink transmission is assumed, where the BS’s antenna array is placed at a height of 25 m. The UT randomly resides in a $120^{\circ}$ cell sector with a radius of 500 m and a guard distance of 35 m to the BS. The number of antennas at the BS and UT is set as $N_{\rm t}=64$ and $N_{\rm r}=16$, respectively. A total of $D=100,000$ channel samples are generated as the training dataset, each with different UT locations to enhance diversity. Two distinct setups are considered, i.e., noiseless and noisy datasets. For the noiseless case (Sections IV-B and IV-C), a simulated dataset =$\mathcal{H}=\{\mathbf{h}^{(i)}\}_{i=1}^{D}$ containing ground truth samples generated via QuaDRiGa is employed; for the noisy case (Section IV-E), the simulated dataset $\mathcal{H}$ serves as a benchmark for deriving normalized LMMSE estimates as given in (32), constituting a noisy dataset ¯=$\bar{\mathcal{H}}=\{\bar{\mathbf{h}}^{(i)}\}_{i=1}^{D}$. Furthermore, a validation dataset containing 100 channel realizations is utilized for hyperparameter tuning.
During training, the number of epochs is set to 500. The batch size is configured as 128. The Adam optimizer with the learning rate fixed as $10^{-4}$ is utilized. For the noise schedule $\{\beta_{t}\}_{t=1}^{T}$ of DM, we employ the linear schedule [^6] We also tried other advanced noise schedules, including the quadratic [50] and exponential schedules [51], but did not observe performance enhancement. prescribed in [53, Table 1] for different numbers of time steps $T$. For the backbone CNN of the denoising network, the maximum channel size is set as $S_{\max}=64$, leading to the channel sizes of different convolutional layers as shown in Fig. 14. Details of the network are illustrated in Appendix A. The gradient scale $s$ of the proposed approach is selected by grid search using the validation dataset. During testing, we generate 100 channel realizations with UT locations that differ from the training dataset to assess the estimation normalized MSE (NMSE) performance defined as
@@
The channel samples are normalized by utilizing the average channel power calculated from the training dataset across all data samples and channel entries, ensuring consistency with the definition of SNR adopted in the system model. The pilot symbols are randomly drawn from the QPSK lattice unless noted otherwise.
@@
Hyperparameters of these baselines are meticulously tuned based on the validation dataset. Note that for the LDAMP and VAE methods, a distinct model is trained for each value of SNR, pilot density $\alpha$, and quantization bit count to maximize their performance due to their supervised learning nature. This results in substantial memory overhead compared to the proposed approach, which adopts a single model that is agnostic to these variables, enabling it to efficiently handle various configurations without the need for retraining. The GAN-based channel estimator from [21] is not selected as a baseline due to its performance saturation at high SNRs, as demonstrated in [25].
@@ IV-B Full-Resolution Channel Estimation
We first evaluate the estimation performance of the proposed DM-based method in the full-resolution receiver. Fig. 3 presents the NMSE as a function of the pilot density $\alpha$ with the SNR fixed as 30 dB. We consider three distinct configurations for the proposed method. The former two are equipped with $T=100$ and $T=1000$, respectively. The last one, also equipped with $T=100$ time steps, utilizes an enhancing strategy to refine the estimate by iterating the update rule 3 rounds at each time index $t\in[0,T/2]$, i.e., the latter half of the reverse process. This strategy, termed “DM T100-Enhanced”, is motivated by the observation that there exists a refinement stage as $t$ approaches 0, during which using additional sampling iterations can significantly enhance the estimation performance [56, 57]. As shown in Fig. 4(a), the “DM T100-Enhanced” variant outperforms the vanilla versions with both $T=100$ and $T=1000$, particularly when $\alpha$ is small. Similar phenomena are noted in the low SNR regime, showcasing the superiority of this variant. Therefore, this variant is maintained for the remaining experiments. Moreover, the figure shows that the NMSE improvement within the range $\alpha\in(0.4,1]$ is relatively gradual compared to the rapid decrease in estimation error within the range $\alpha\in[0.2,0.4]$. This result reveals that the prior incorporated in the DM can provide sufficient information for approaching the recovery performance achieved in a fully determined inverse problem.
@@
Fig. 4 compares the NMSE and latency of the proposed method with baselines when the pilot density is $\alpha=1$. As shown in Fig. 4(a), the proposed DM-based approach outperforms all the compared schemes by a wide margin in estimation accuracy except for the high-complexity SGM. Due to the absence of true covariance information, the LMMSE estimator only uses an estimated sample covariance, hence lagging behind the DM that can capture deep-level propagation features. In addition, we compare the proposed approach with the DM-based channel estimator introduced in [26], which fails to achieve the same performance as the proposed approach. This disparity can be attributed to the fact that the theoretical framework and algorithm design presented in [26] rely on the assumption of orthogonal pilot sequences and full pilot measurements. In contrast, the proposed method is not restricted by this assumption, enabling its applicability to scenarios with non-orthogonal pilot sequences, under-determined measurements (i.e., $N_{\rm p}<N_{\rm t}$), and low-resolution ADCs. Given the limitations of the DM-based approach in [26] within these contexts, we have opted not to include it as a baseline for comparison in the subsequent simulations. Fig. 4(b) presents the online estimation latency per channel realization across the compared methods. The latency is evaluated on a machine with an NVIDIA Tesla V100 GPU and an Intel Xeon E5-4627 CPU. The figure demonstrates that the proposed method achieves notable latency reduction compared to the SGM, making it one of the most efficient methods among those compared. This advantage can be attributed to the lightweight network architecture, substantially reduced time steps, and low computational overhead at each step. Considering the comparable performance to SGM as depicted in Fig. 4(a), our method can be viewed as a favorable option for wireless channel estimation featuring real-time implementation.
@@
To verify the robustness of the proposed deep generative prior-based method, we consider a mismatched channel setting in Fig. 7. Specifically, we deploy the DM trained using LOS channel samples in a non-line-of-sight (NLOS) scenario (with other configurations remaining unchanged) for channel estimation without any retraining or adaptation. We also present the NMSE of benchmark models trained with NLOS channel samples. As shown in the figure, the performance loss of the proposed method in mismatched setups is very small compared to that of the LMMSE and LDAMP, similar to the SGM baseline. This can be attributed to the powerful distribution coverage capability of generative DMs [17].
@@ IV-C Low-Resolution Channel Estimation
Figure 10: Performance with various ADC resolutions when $\alpha=0.6$.
@@
Figure 11: Block diagram of the considered system setup for evaluating the end-to-end BER performance.
Figure 12: End-to-end BER performance of the downlink communication link using the proposed estimator.
IV-D End-to-end Bit Error Rate (BER) Performance
In this subsection, we evaluate the proposed estimator using the end-to-end BER performance within a downlink transmission system (Fig. 11) as the metric. The channel setup and the antenna configuration align with the description provided in Sec. IV-A. For pilot transmissions, a pilot matrix $\mathbf{P}\in\mathbb{C}^{N_{\rm t}\times N_{\rm p}}$ with $N_{\rm p}=38$ is sent by the transmitter, and the received signal $\mathbf{Y}$ at the user is utilized for channel estimation, resulting in $\hat{\mathbf{H}}$. During the data transmission phase, a rate-1/2 low-density parity-check (LDPC) code with a codeword size of 648 bits is adopted to encode the information bits $\mathbf{b}$, followed by quadrature amplitude modulation (QAM) to generate the data symbols. The modulation scheme is selected as 64-QAM. The data symbols are partitioned into $N_{\rm s}=8$ streams, denoted as a data matrix $\mathbf{D}$ over multiple channel realizations, and are subsequently precoded for transmission over $N_{\rm t}=64$ antennas. We assume perfect channel state information (CSI) at the transmitter, and the digital precoder $\mathbf{V}$ is constructed by the first $N_{\rm s}$ right-singular vectors of the channel matrix $\mathbf{H}$, obtained through singular value decomposition (SVD). At the receiver, the data observation $\mathbf{Y}_{\rm D}$ is LMMSE equalized given the estimated channel $\hat{\mathbf{H}}$, soft demapped, and decoded by the LDPC decoder to produce the estimated information bits $\hat{\mathbf{b}}$, with BER computed after $5\times 10^{4}$ codewords are transmitted.
Fig. 12 compares the BER performance when different channel estimators are used. The LASSO, EM-GM-AMP, and VAE estimators are not considered in this comparison due to their estimation NMSE floors at high SNRs, which hinders effective data transmission. The end-to-end BER performance with perfect CSI at the receiver is adopted as a performance upper bound for comparison. The figure illustrates that our proposed estimator achieves a significant reduction in BER at high SNRs, outperforming the LMMSE- and LDAMP-based approaches by more than 5 dB and surpassing the more complicated SGM-based approach. These findings underscore the significance of the high-fidelity channel acquisition provided by the proposed DM-based estimator in improving end-to-end system performance.
IV-E Effectiveness of SURE-DM
@@ V Conclusion
Although the proposed method offers significant advantages, further endeavors are needed to address the theoretical analysis on its robustness to out-of-distribution channel data, as well as the fast adaptation under limited data samples. Additionally, theoretical justifications for the heuristic enhancements used in the proposed approach, such as iterative refinement and the gradient scaling parameter, warrant further investigation.